Determine if the following pairs of functions are inverses of each other; you must show your work that proves your answer:
a. f(x) = 4x and g(x) = 0.25x
b. f(x) = x – 4 and g(x) = 2x + 4

Answer by eperette(173)

Concept you should know:
F(x) and G(x) are inverses of each other if
F(G(x))=G(F(x))=1x
********************************************************
a. f(x) = 4x and g(x) = 0.25x
f(g(x))=f(0.25x)=4(0.25x)=x
g(f(x))=g(4x)=0.25(4x)=x
they are inverses of each other
*********************************************************
b. f(x) = x – 4 and g(x) = 2x + 4
f(g(x))=f(2x+4)= 2x + 4 -4 =2x
g(f(x))=g(x-4) = 2(x-4) + 4 = 2x - 8 + 4 = 2x -4
they are not inverses of each other